# A Neural Surrogate-Enhanced Multi-Method Framework for Robust Wing Design Optimization

### v1

@@ Abstract
This paper introduces a modular and scalable design optimization framework for the wing design process that enables faster early-phase design while ensuring aerodynamic stability. The pipeline starts with the generation of initial wing geometries and then proceeds to optimize the wing using several algorithms. Aerodynamic performance is assessed using a Vortex Lattice Method (VLM) applied to a carefully selected dataset of wing configurations. These results are employed to develop surrogate neural network models, which can predict lift and drag rapidly and accurately. The stability evaluation is implemented by setting the control surfaces and components to fixed positions in order to have realistic flight dynamics. The approach unifies and compares several optimization techniques, including Particle Swarm Optimization (PSO), Genetic Algorithms (GA), gradient-based MultiStart methods, Bayesian optimization, and Lipschitz optimization. Each method ensures constraint management via adaptive strategies and penalty functions, where the targets for lift and design feasibility are enforced. The progression of aerodynamic characteristics and geometries over the optimization iterations will be investigated in order to clarify each algorithm’s convergence characteristics and performance efficiency. Our results show improvement in aerodynamic qualities and robust stability properties, offering a mechanism for wing design at speed and precision. In the interest of reproducibility and community development, the complete implementation is publicly available at https://github.com/AmirHosseinGhaemi2000/CHIMERA.
@@ 1.1 Background and Related Work
Adjoint-based approaches continue to play a significant role in aerodynamic shape optimization, as these approaches can reliably compute the gradient with respect to many design variables at a relatively low cost, and recent research has implemented stability constraints in the optimization process [1]. To reduce the reliance on adjoint solvers, adjoint-free methods have been developed for complex geometries and configurations, such as the Common Research Model wing, that achieved similar fidelity without having to solve adjoint equations [2]. Surrogate-assisted adjoint techniques have also been explored, for example, by integrating radial basis functions such that they converge more effectively [3]. Gradient prediction techniques based on dynamic-mode decomposition have lessened the computational time for gradient solutions by more than 80%, and thus accelerated the computation [4]. Unified adjoint formulations for lattice Boltzmann methods have broadened the applicability to viscous/inviscid, compressible/incompressible, and 2D/3D flows [5]. Robust design regulations have been integrated into adjoint frameworks by coupling uncertainty quantification via polynomial chaos expansion, as a way to obtain designs that optimized aerodynamic performance while yielding robustness under multiple uncertainty sources [6].
@@
In parallel, deep learning and surrogate modeling are providing new alternatives to aerodynamic design. Neural-network surrogates that are integrated with gradient-based optimization improve convergence speed and the quality of the solutions [7], while data-driven techniques provide efficient exploration of large wing design spaces [8]. The applications are extending from turbo-machinery [9], generative modeling of unconventional and non-standard configurations [10, 11], and hybrid methods integrating genetic algorithms and deep neural networks for airfoil optimization [12]. Adaptive machine learning-based methods have allowed for optimization of finite wings for different sets of flight conditions [13], whereas deep learning techniques have managed to outperform reduced-order methods in more complex non-linear transonic regimes [14]. Geometric filtering methods provide further constraints on design spaces on physically valid yet aerodynamically innovative shapes [15].
@@
Multi-fidelity and transfer-learning approaches have been extensively implemented for balancing accuracy and computational efficiency. For instance, CNN-based multi-fidelity surrogates [16, 17], surrogate-assisted optimization of aircraft wings [18], and transfer-learning frameworks aimed at the transfer of aerodynamic knowledge across modeling problems [19]. The inverse design approaches that combine multi-fidelity simulations with machine learning have extended the applicability as well [20]. Component-level studies, such as minimizing wing–fuselage interference, further highlight the value of multi-fidelity and hybrid optimization approaches [21].
@@
There has been a growing focus on multi-objective optimization, uncertainty quantification, and robustness. New contributions include efficient handling of geometric constraints [22], improved global search strategies for expensive problems [23], multi-objective particle swarm optimization methods tailored to turbo-machinery [24], and the use of deep-learning-aided evolutionary algorithms to produce the Pareto front [25]. Many robust optimization approaches contain uncertainty quantification by implementing radial basis function (RBF) surrogates [26], stochastic gradient-based optimization [27], as well as uncertainty-aware surrogates for complex aerodynamic problems. Off-design performance analysis is derived from imposed constraints that accurately emulate real-world flight environments while requiring very few simulations [28].
@@
In addition to conventional aircraft, novel optimization methods have also been implemented for unconventional and biologically inspired configurations. The triangular airfoils for Martian flight [29], corrugated wing designs refined with reinforcement learning [30], Unmanned Aerial Vehicle (UAV) penetration strategies with six degrees of freedom using deep reinforcement learning [31], and aero-structural systems that balance weight and stiffness for high-aspect-ratio wings [32], are all few of the studies implementing novel optimization methods. Hybrid surrogate–metaheuristic methods were also employed in lightweight structural designs for flexible wings [33].
@@
There are still persisting fundamental difficulties with respect to sensitivity to initial conditions, representation of geometry, and robustness to uncertainty. There has been research around conducting statistical analysis of optimization robustness [34], dynamically controlling geometry to maintain quality of the mesh [35], compact modal design variable parameterizations [36], and optimization using a large database of design variables [37]. Distributionally robust formulations have further improved resiliency to uncertain data [38]. Reviews have detailed the continuing challenges in balancing fidelity, cost, and multidisciplinary complexity [39].
@@
Lastly, recent works have associated mission-level objectives with aerodynamic design. [40] Reinforcement learning with long short-term memories (LSTMs) was used in UAVs, in order to optimize communication under energy constraints [40]. In the configuration level, machine learning-based optimization has also helped accelerate the design of unconventional configurations such as the box wings, demonstrating meaningful efficiency gains that were validated with Reynolds-Averaged Navier–Stokes equations (RANS) [41]. Overall, these advancements suggest a shift towards combining surrogates, physics-based models, and intelligent algorithms into a unified, scalable framework.
@@ 4.1 Implementation Details
Aerodynamic data were created using a vortex-lattice based solver [42] and stability derivatives were calculated by linearizing the equations of motion about trimmed flight states (derivations shown in A). The input design variables included eight wing and flight parameters: root chord, angle of attack, sweep angle, half-span, twist angle, taper ratio, dihedral angle, and free-stream velocity. Their ranges, listed in Table 1, were selected to allow enough coverage of realistic ultralight aircraft configurations.
@@ Appendix A Equations of Motion and Stability
This appendix summarizes the kinematics, aerodynamic load formulation, rigid-body dynamics, and the linear stability derivatives employed in the analysis of fixed-wing aircraft. All formulas in this appendix are drawn from [43]. Conventions: body axes $(x,y,z)$ point forward, right, and down; angular rates $(p,q,r)$ are positive about $(x,y,z)$; gravity $g>0$ acts downward. Moments are taken about the center of gravity (CG). Aerodynamic coefficients are referenced to $(S,b,c)$; unless noted, the aerodynamic moment reference is at the quarter-chord.
@@ Appendix B Optimization Methods, Hyperparameters, & Pseudo Codes
where (4) is minimized by the gradient-based method with the equality constraint (3), while (5) is used by the derivative-free global methods [44, 45]. All methods keep $\mathbf{lb}\leq\mathbf{x}\leq\mathbf{ub}$ at every step.
@@ B.1 Gradient-based Method (Multi-start + Interior points)
The necessary optimality conditions at a local solution $\mathbf{x}^{\star}$ are the Karush–Kuhn–Tucker (KKT) conditions [46, 47, 48]. Let $\lambda\in\mathbb{R}$ be the multiplier of (3), and let $\boldsymbol{\alpha},\boldsymbol{\beta}\in\mathbb{R}^{8}_{\geq 0}$ be the multipliers for the lower and upper bounds, respectively. Define the Lagrangian
@@ B.2 Particle Swarm Optimization (PSO)
where $W$ is the inertia, $c_{1},c_{2}$ are cognitive/social weights, $\mathbf{u}_{1},\mathbf{u}_{2}\sim\mathcal{U}[0,1]^{8}$ act componentwise, $\mathbf{p}$ is the particle best, and $\mathbf{g}$ is the swarm best. PSO minimizes the penalized objective $f_{\text{p}}$ in (5). Classical PSO and its convergence analysis follow [49, 50]. The full PSO pipeline and hyperparameters appear in Tables 14 and 15, respectively.
@@ B.3 Genetic Algorithm (GA)
The GA operates on a population using fitness proportional to $f_{\text{p}}$. Parents are selected (e.g., stochastic uniform), recombined by convex or heuristic crossover, and mutated by random perturbations, with each offspring projected back to the box $\mathbf{lb},\mathbf{ub}]$. This method minimizes the penalized objective $f_{\text{p}}$ in (5). Foundational references include Holland [51] and Goldberg [52]. The end-to-end pipeline and hyperparameters are shown in Tables 16 and 17.
@@ B.4 Bayesian Optimization
where $\mu_{t}(\mathbf{x})$ and $\sigma_{t}(\mathbf{x})$ are the surrogate posterior mean and standard deviation, and $f_{\min}$ is the incumbent objective value. The pipeline and hyperparameters are in Tables 18 and 19. This lineage follows Kushner [53] and Mockus et al. [54], with the modern EGO formulation of Jones et al. [55].
@@ B.5 Lipschitz Global Search with Local Constrained Refinement
with reliability factor $r>1$. When a selected box is sufficiently small, a local refinement solves the constrained problem (6) starting at the box center. The pipeline and hyperparameters are shown in Tables 20 and 21. See [56, 57, 58] for foundational Lipschitz global optimization.

### v2

@@ Abstract
This paper introduces a modular and scalable design optimization framework for the wing design process that enables faster early-phase design while ensuring aerodynamic stability. The pipeline starts with the generation of initial wing geometries and then proceeds to optimize the wing using several algorithms. Aerodynamic performance is assessed using a Vortex Lattice Method (VLM) applied to a carefully selected dataset of wing configurations. These results are employed to develop surrogate neural network models, which can predict lift and drag rapidly and accurately. The stability evaluation is implemented by setting the control surfaces and components to fixed positions in order to have realistic flight dynamics. The approach unifies and compares several optimization techniques, including Particle Swarm Optimization (PSO), Genetic Algorithms (GA), gradient-based MultiStart methods, Bayesian optimization, and Lipschitz optimization. Each method ensures constraint management via adaptive strategies and penalty functions, where the targets for lift and design feasibility are enforced. The progression of aerodynamic characteristics and geometries over the optimization iterations will be investigated in order to clarify each algorithm’s convergence characteristics and performance efficiency. Our results show improvement in aerodynamic qualities and robust stability properties, offering a mechanism for wing design at speed and precision. In the interest of reproducibility and community development, the complete implementation is publicly available on GitHub.
@@ 1.1 Background and Related Work
Adjoint-based approaches continue to play a significant role in aerodynamic shape optimization, as these approaches can reliably compute the gradient with respect to many design variables at a relatively low cost, and recent research has implemented stability constraints in the optimization process [he2025adjoint]. To reduce the reliance on adjoint solvers, adjoint-free methods have been developed for complex geometries and configurations, such as the Common Research Model wing, that achieved similar fidelity without having to solve adjoint equations [li2021adjoint]. Surrogate-assisted adjoint techniques have also been explored, for example, by integrating radial basis functions such that they converge more effectively [abergo2023aerodynamic]. Gradient prediction techniques based on dynamic-mode decomposition have lessened the computational time for gradient solutions by more than 80%, and thus accelerated the computation [chen2024dynamic]. Unified adjoint formulations for lattice Boltzmann methods have broadened the applicability to viscous/inviscid, compressible/incompressible, and 2D/3D flows [khouzani2025mathematical]. Robust design regulations have been integrated into adjoint frameworks by coupling uncertainty quantification via polynomial chaos expansion, as a way to obtain designs that optimized aerodynamic performance while yielding robustness under multiple uncertainty sources [shi2023robust].
@@
In parallel, deep learning and surrogate modeling are providing new alternatives to aerodynamic design. Neural-network surrogates that are integrated with gradient-based optimization improve convergence speed and the quality of the solutions [wu2024gradient], while data-driven techniques provide efficient exploration of large wing design spaces [hasan2025aerodynamic]. The applications are extending from turbo-machinery [esfahanian2024aerodynamic], generative modeling of unconventional and non-standard configurations [jabon2024aerodynamic, wang2023airfoil], and hybrid methods integrating genetic algorithms and deep neural networks for airfoil optimization [wu2023airfoil]. Adaptive machine learning-based methods have allowed for optimization of finite wings for different sets of flight conditions [zhang2024adaptive], whereas deep learning techniques have managed to outperform reduced-order methods in more complex non-linear transonic regimes [sabater2022fast]. Geometric filtering methods provide further constraints on design spaces on physically valid yet aerodynamically innovative shapes [li2021deep].
@@
Multi-fidelity and transfer-learning approaches have been extensively implemented for balancing accuracy and computational efficiency. For instance, CNN-based multi-fidelity surrogates [tao2024multi, du2025aerodynamic], surrogate-assisted optimization of aircraft wings [nikolaou2025multi], and transfer-learning frameworks aimed at the transfer of aerodynamic knowledge across modeling problems [wu2024efficient]. The inverse design approaches that combine multi-fidelity simulations with machine learning have extended the applicability as well [grbcic2024efficient]. Component-level studies, such as minimizing wing–fuselage interference, further highlight the value of multi-fidelity and hybrid optimization approaches [matos2025aerodynamic].
@@
There has been a growing focus on multi-objective optimization, uncertainty quantification, and robustness. New contributions include efficient handling of geometric constraints [wang2023efficient], improved global search strategies for expensive problems [deng2022exploitation], multi-objective particle swarm optimization methods tailored to turbo-machinery [li2025optimization], and the use of deep-learning-aided evolutionary algorithms to produce the Pareto front [bakar2022multi]. Many robust optimization approaches contain uncertainty quantification by implementing radial basis function (RBF) surrogates [asouti2023radial], stochastic gradient-based optimization [jofre2022rapid], as well as uncertainty-aware surrogates for complex aerodynamic problems. Off-design performance analysis is derived from imposed constraints that accurately emulate real-world flight environments while requiring very few simulations [li2023efficient].
@@
In addition to conventional aircraft, novel optimization methods have also been implemented for unconventional and biologically inspired configurations. The triangular airfoils for Martian flight [caros2023optimization], corrugated wing designs refined with reinforcement learning [hirooka2024shape], Unmanned Aerial Vehicle (UAV) penetration strategies with six degrees of freedom using deep reinforcement learning [zhuang2024optimization], and aero-structural systems that balance weight and stiffness for high-aspect-ratio wings [luo2024aero], are all few of the studies implementing novel optimization methods. Hybrid surrogate–metaheuristic methods were also employed in lightweight structural designs for flexible wings [li2024hybrid].
@@
There are still persisting fundamental difficulties with respect to sensitivity to initial conditions, representation of geometry, and robustness to uncertainty. There has been research around conducting statistical analysis of optimization robustness [streuber2021evaluating], dynamically controlling geometry to maintain quality of the mesh [streuber2023improved], compact modal design variable parameterizations [kedward2023generic], and optimization using a large database of design variables [li2021data]. Distributionally robust formulations have further improved resiliency to uncertain data [chen2024data]. Reviews have detailed the continuing challenges in balancing fidelity, cost, and multidisciplinary complexity [martins2022aerodynamic].
@@
Lastly, recent works have associated mission-level objectives with aerodynamic design. [he2024deep] Reinforcement learning with long short-term memories (LSTMs) was used in UAVs, in order to optimize communication under energy constraints [he2024deep]. In the configuration level, machine learning-based optimization has also helped accelerate the design of unconventional configurations such as the box wings, demonstrating meaningful efficiency gains that were validated with Reynolds-Averaged Navier–Stokes equations (RANS) [hasan2025enhancing]. Overall, these advancements suggest a shift towards combining surrogates, physics-based models, and intelligent algorithms into a unified, scalable framework.
@@ 4.1 Implementation Details
Aerodynamic data were created using a vortex-lattice based solver [melin2000vortex] and stability derivatives were calculated by linearizing the equations of motion about trimmed flight states (derivations shown in A). The input design variables included eight wing and flight parameters: root chord, angle of attack, sweep angle, half-span, twist angle, taper ratio, dihedral angle, and free-stream velocity. Their ranges, listed in Table 1, were selected to allow enough coverage of realistic ultralight aircraft configurations.
@@ Appendix A Equations of Motion and Stability
This appendix summarizes the kinematics, aerodynamic load formulation, rigid-body dynamics, and the linear stability derivatives employed in the analysis of fixed-wing aircraft. All formulas in this appendix are drawn from [roskam1997airplane]. Conventions: body axes $(x,y,z)$ point forward, right, and down; angular rates $(p,q,r)$ are positive about $(x,y,z)$; gravity $g>0$ acts downward. Moments are taken about the center of gravity (CG). Aerodynamic coefficients are referenced to $(S,b,c)$; unless noted, the aerodynamic moment reference is at the quarter-chord.
@@ Appendix B Optimization Methods, Hyperparameters, & Pseudo Codes
where (4) is minimized by the gradient-based method with the equality constraint (3), while (5) is used by the derivative-free global methods [Bertsekas1982, Forrester2008]. All methods keep $\mathbf{lb}\leq\mathbf{x}\leq\mathbf{ub}$ at every step.
@@ B.1 Gradient-based Method (Multi-start + Interior points)
The necessary optimality conditions at a local solution $\mathbf{x}^{\star}$ are the Karush–Kuhn–Tucker (KKT) conditions [Karush1939, Kuhn1951, Nocedal2006]. Let $\lambda\in\mathbb{R}$ be the multiplier of (3), and let $\boldsymbol{\alpha},\boldsymbol{\beta}\in\mathbb{R}^{8}_{\geq 0}$ be the multipliers for the lower and upper bounds, respectively. Define the Lagrangian
@@ B.2 Particle Swarm Optimization (PSO)
where $W$ is the inertia, $c_{1},c_{2}$ are cognitive/social weights, $\mathbf{u}_{1},\mathbf{u}_{2}\sim\mathcal{U}[0,1]^{8}$ act componentwise, $\mathbf{p}$ is the particle best, and $\mathbf{g}$ is the swarm best. PSO minimizes the penalized objective $f_{\text{p}}$ in (5). Classical PSO and its convergence analysis follow [Kennedy1995, Clerc2002]. The full PSO pipeline and hyperparameters appear in Tables 14 and 15, respectively.
@@ B.3 Genetic Algorithm (GA)
The GA operates on a population using fitness proportional to $f_{\text{p}}$. Parents are selected (e.g., stochastic uniform), recombined by convex or heuristic crossover, and mutated by random perturbations, with each offspring projected back to the box $\mathbf{lb},\mathbf{ub}]$. This method minimizes the penalized objective $f_{\text{p}}$ in (5). Foundational references include Holland [Holland1975] and Goldberg [Goldberg1989]. The end-to-end pipeline and hyperparameters are shown in Tables 16 and 17.
@@ B.4 Bayesian Optimization
where $\mu_{t}(\mathbf{x})$ and $\sigma_{t}(\mathbf{x})$ are the surrogate posterior mean and standard deviation, and $f_{\min}$ is the incumbent objective value. The pipeline and hyperparameters are in Tables 18 and 19. This lineage follows Kushner [Kushner1964] and Mockus et al. [Mockus1978], with the modern EGO formulation of Jones et al. [Jones1998].
@@ B.5 Lipschitz Global Search with Local Constrained Refinement
with reliability factor $r>1$. When a selected box is sufficiently small, a local refinement solves the constrained problem (6) starting at the box center. The pipeline and hyperparameters are shown in Tables 20 and 21. See [Piyavskii1972, Shubert1972, Hansen1992] for foundational Lipschitz global optimization.